OpenStudy (anonymous):

Solve the absolute value equation for x. |2x−4|=12

OpenStudy (anonymous):

x= -8, ?

OpenStudy (anonymous):

x=-8

myininaya (myininaya):

You have two equations to solve. 2x-4=12 or 2x-4=-12

myininaya (myininaya):

You already solved the first equation. Just solve the other one 2x-4=-12

OpenStudy (anonymous):

which is 8 rite?

myininaya (myininaya):

nope

myininaya (myininaya):

2x-4=-12 How do we get x by itself? First what action must we perform in order to get 2x by itself?

OpenStudy (anonymous):

-4, thank you!

OpenStudy (anonymous):

move it to the other side

myininaya (myininaya):

gj @jdz359